# Llamba: Scaling Distilled Recurrent Models for Efficient Language Processing

### v1

@@ 1 Introduction
Table 1: Comparison of downstream performance (accuracy %) in zero-shot settings across various models. We use the following abbreviations: ARC Challenge (ARC-C), ARC Easy (ARC-E), Physical Interaction QA (PIQA), Winogrande (WG), HellaSwag (HS), Lambada (LMB), Massive Multitask Language Understanding (MMLU), and OpenBookQA (OBQA). For ARC-C, ARC-E, PIQA, HellaSwag, and OBQA, we use normalized logits’ results. Averages are computed over all 8 tasks. Along with accuracy, each model is annotated with the number of training or distillation tokens (in trillions) and its architecture—Recurrent (R), Transformer (T), or Hybrid (H). For models with a sliding window, the window size is also specified. This table shows that all Llamba models achieve competitive performance against both Transformer-based and SSM-based models, despite their recurrent layers being trained on significantly fewer tokens. See Tables 3 and 4 for the full few-shot comparisons.
@@ 2 Related Work
Various methods have been proposed to distill large Transformer-based models into more efficient architectures while maintaining performance. LoLCATs (Zhang et al., 2024) introduces a linearization approach that replaces softmax attention with linear attention through attention transfer, followed by low-rank adaptation, enabling the creation of large-scale linearized models with improved efficiency. (Wang et al., 2025) leverages the State-Space Duality (SSD) in Dao and Gu (2024) to transfer the linear projection weights from the attention layers into Mamba-based models. Their experiments include hybrid models with an increasing proportion of interleaved attention layers, demonstrating that a balanced combination of state-space models (SSMs) and attention preserves performance while improving efficiency. MOHAWK (Bick et al., 2024) distills Transformers into SSMs through progressive alignment, allowing subquadratic models to leverage Transformer training resources effectively. These approaches demonstrate the viability of distilling computationally expensive Transformers into efficient models while retaining competitive performance.
@@ Appendix A Comparison of Downstream Performance
We evaluate a range of language models across multiple benchmarks, measuring their performance in both zero-shot and few-shot settings. ARC-Challenge and ARC-Easy [Clark et al., 2018], and PIQA [Bisk et al., 2019] were evaluated with 0 and 25 shots, Winogrande [Sakaguchi et al., 2019] with 0 and 5 shots, HellaSwag [Zellers et al., 2019] and OpenBookQA Mihaylov et al. [2018] with 0 and 10 shots, and Lambada [Paperno et al., 2016] and MMLU [Hendrycks et al., 2021] with 0 and 5 shots. The results, reported in Tables 3 and 4, use normalized logits for ARC-Challenge, ARC-Easy, PIQA, HellaSwag, and OpenBookQA. These benchmarks cover a range of reasoning, commonsense, and language comprehension tasks, providing insight into the models’ ability to process different types of contextual dependencies.
The evaluation includes both quadratic and sub-quadratic architectures. Specifically, we evaluate the Transformer-based Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B [Touvron et al., 2023], as well as Qwen2.5-3B [Yang et al., 2024]. Falcon-Mamba-7B and Falcon3-Mamba-7B [Zuo et al., 2024], RecurrentGemma-2B and RecurrentGemma-9B [Botev et al., 2024], and Mamba2-2.8B [Dao and Gu, 2024] are included as models integrating recurrent components. Additionally, we evaluate Zamba2-7B [Glorioso et al., 2024a], a hybrid that incorporates both attention and SSMs. To provide a more comprehensive comparison, we also assess Llamba-1B, Llamba-3B, and Llamba-8B, which apply recurrent architectures across different scales. These models vary in their training strategies, data sources, and architectural choices, allowing for a broad comparison of language model capabilities across different methodological approaches.
Table 3: Comparison of downstream performance (accuracy %) across various models in zero-shot and few-shot settings. For ARC-Challenge, ARC-Easy, and PIQA, we have used normalized logits’ results. Along with accuracy, each model is annotated with the number of training or distillation tokens (in trillions) and its architecture—Recurrent (R), Transformer (T), or Hybrid (H). For models with a sliding window, the window size is also specified.
@@
Table 4: Comparison of downstream performance (accuracy %) across various models in zero-shot and few-shot settings. For HellaSwag and OpenBookQA, we have used normalized logits’ results. Along with accuracy, each model is annotated with the number of training or distillation tokens (in trillions) and its architecture—Recurrent (R), Transformer (T), or Hybrid (H). For models with a sliding window, the window size is also specified.

### v2

@@ 1 Introduction
Table 1: Comparison of downstream performance (accuracy %) in zero-shot settings across various models. We use the following abbreviations: ARC Challenge (ARC-C), ARC Easy (ARC-E), Physical Interaction QA (PIQA), Winogrande (WG), HellaSwag (HS), Lambada (LMB), Massive Multitask Language Understanding (MMLU), and OpenBookQA (OBQA). For ARC-C, ARC-E, PIQA, HellaSwag, and OBQA, we use normalized logits’ results. Averages are computed over all 8 tasks. Along with accuracy, each model is annotated with the number of training or distillation tokens (in trillions) and its architecture—Recurrent (R), Transformer (T), or Hybrid (H). For models with a sliding window, the window size is also specified. This table shows that all Llamba models achieve competitive performance against both Transformer-based and SSM-based models, despite their recurrent layers being trained on significantly fewer tokens. We use an instruct-tuned version whenever one is available; however, we exclude this label for brevity. See Tables 4 and 5 for the full few-shot comparisons.
@@ 2 Related Work
Various methods have been proposed to distill large Transformer-based models into more efficient architectures while maintaining performance. SUPRA (Mercat et al., 2024) propose a procedure to linearize softmax attention into a form of linear attention by copying the weight matrices and fine-tuning. LoLCATs (Zhang et al., 2024) introduces a linearization approach that replaces softmax attention with linear attention through attention transfer, followed by low-rank adaptation, enabling the creation of large-scale linearized models with improved efficiency. (Wang et al., 2025) leverages the State-Space Duality (SSD) in Dao and Gu (2024) to transfer the linear projection weights from the attention layers into Mamba-based models. Their experiments include hybrid models with an increasing proportion of interleaved attention layers, demonstrating that a balanced combination of state-space models (SSMs) and attention preserves performance while improving efficiency. MOHAWK (Bick et al., 2024) distills Transformers into SSMs through progressive alignment, allowing subquadratic models to leverage Transformer training resources effectively. These approaches demonstrate the viability of distilling computationally expensive Transformers into efficient models while retaining competitive performance.
@@ 6.1 Performance
Comparison Against Distilled Models.
Table 3 compares Llamba-8B with other distilled models of similar size. We specifically focus on MMLU, which is known to be difficult for recurrent models (Waleffe et al., 2024), and has the biggest gap for distilled models from prior work. Llamba significantly improves MMLU relative to the teacher model.
We found that MMLU performance takes much longer to improve compared to other benchmarks in our end-to-end distillation. Llamba reached teacher performance on other tasks in a very small number of tokens, while MMLU took longer to improve.
We also note that some of the baselines are actually hybrid models, which have a 1:1 ratio of attention to recurrent layers. We note that even sliding window attention has a strong effect because the MMLU context size is very small. Although Llamba still has a gap to the teacher model, we consider this result a large step forward for the performance of distilled recurrent models.
Model Arch. Tokens (B) Teacher MMLU Score Relative Score (%)
SUPRA R 100 Mistral 7B 34.2 24.6
Mamba2-Llama 3 R 20 Llama 3 8B 43.2 43.8
Mamba2-Llama 3 H 20 Llama 3 8B 56.7 76.2
LoLCATs H 0.04 Mistral 7B 51.4 70.5
LoLCATs H 0.04 Llama 3 8B 52.8 66.8
Llamba-8B R 12 Llama 3.1 8B 60.0 80.6
Table 3: MMLU performance of different models along with their respective architectures, training tokens, and teachers. Along with accuracy, each model is annotated with the number of training or distillation tokens (in trillions) and its architecture—Recurrent (R) or Hybrid (H). The relative score indicates the MMLU score relative to the teacher’s score where the lower threshold is taken to be 25 (random guessing).
@@ Appendix A Comparison of Downstream Performance
We evaluate a range of language models across multiple benchmarks, measuring their performance in both zero-shot and few-shot settings. ARC-Challenge and ARC-Easy [Clark et al., 2018], and PIQA [Bisk et al., 2019] were evaluated with 0 and 25 shots, Winogrande [Sakaguchi et al., 2019] with 0 and 5 shots, HellaSwag [Zellers et al., 2019] and OpenBookQA Mihaylov et al. [2018] with 0 and 10 shots, and Lambada [Paperno et al., 2016] and MMLU [Hendrycks et al., 2021] with 0 and 5 shots. The results, reported in Tables 4 and 5, use normalized logits for ARC-Challenge, ARC-Easy, PIQA, HellaSwag, and OpenBookQA. These benchmarks cover a range of reasoning, commonsense, and language comprehension tasks, providing insight into the models’ ability to process different types of contextual dependencies.
Along with Llamba-1B, Llamba-3B, and Llamba-8B, we evaluate different quadratic and sub-quadratic architectures. Specifically, we include Transformer-based models such as Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, and Llama-3.1-8B-Instruct [Touvron et al., 2023], as well as Qwen2.5-3B [Yang et al., 2024]. Models integrating recurrent components include Falcon-Mamba-7B-Instruct and Falcon3-Mamba-7B-Instruct [Zuo et al., 2024], RecurrentGemma-2B-it and RecurrentGemma-9B-it [Botev et al., 2024], and Mamba varoatopms [Gu and Dao, 2023, Dao and Gu, 2024]. Additionally, we evaluate Zamba2-7B-Instruct [Glorioso et al., 2024a], a hybrid incorporating both attention and SSMs. Whenever an instruct-tuned version is available, we use it, as these models generally perform better. These models differ in training strategies, data sources, and architectural choices, enabling a broad comparison of language model capabilities across different methodologies.
Table 4: Comparison of downstream performance (accuracy %) across various models in zero-shot and few-shot settings. For ARC-Challenge, ARC-Easy, and PIQA, we have used normalized logits’ results. Along with accuracy, each model is annotated with the number of training or distillation tokens (in trillions) and its architecture—Recurrent (R), Transformer (T), or Hybrid (H). For models with a sliding window, the window size is also specified. We use an instruct-tuned version whenever one is available; however, we exclude this label for brevity.
@@
Table 5: Comparison of downstream performance (accuracy %) across various models in zero-shot and few-shot settings. For HellaSwag and OpenBookQA, we have used normalized logits’ results. Along with accuracy, each model is annotated with the number of training or distillation tokens (in trillions) and its architecture—Recurrent (R), Transformer (T), or Hybrid (H). For models with a sliding window, the window size is also specified. We use an instruct-tuned version whenever one is available; however, we exclude this label for brevity.